Check if matrix A=

1 a b

0 1 c

0 0 1

Where a,b,c∈R

is an abelian group with respect to matrix multiplication.


1
Expert's answer
2021-06-08T08:48:01-0400

Let us check whether the set "A=\Big\{ \begin{pmatrix}1 & a & b\\ 0 & 1 & c\\ 0 & 0 & 1\end{pmatrix}:a,b,c\in\mathbb R\Big\}" is an Abelian group with respect to matrix multiplication. It is sufficiently to prove that "A" is a subgroup of the general linear group "GL_3(\mathbb R)."


Let "\begin{pmatrix}1 & a_1 & b_1\\ 0 & 1 & c_1\\ 0 & 0 & 1\end{pmatrix}, \begin{pmatrix}1 & a_2 & b_2\\ 0 & 1 & c_2\\ 0 & 0 & 1\end{pmatrix}\in A."


Taking into account that


"\begin{pmatrix}1 & a_1 & b_1\\ 0 & 1 & c_1\\ 0 & 0 & 1\end{pmatrix}\cdot \begin{pmatrix}1 & a_2 & b_2\\ 0 & 1 & c_2\\ 0 & 0 & 1\end{pmatrix}=
\begin{pmatrix}1 & a_2+a_1 & b_2+a_1c_2+b_1\\ 0 & 1 & c_2+c_1\\ 0 & 0 & 1\end{pmatrix}\in A" and


"\begin{pmatrix}1 & a & b\\ 0 & 1 & c\\ 0 & 0 & 1\end{pmatrix}^{-1}=
\begin{pmatrix}1 & -a & ac-b\\ 0 & 1 & -c\\ 0 & 0 & 1\end{pmatrix}\in A," we conclude that "A" is a subgroup of "GL_3(\mathbb R),"

that is "A" is a group with respect to matrix multiplication.


Since "\begin{pmatrix}1 & 1 & 0\\ 0 & 1 & 0\\ 0 & 0 & 1\end{pmatrix}\cdot \begin{pmatrix}1 & 0 & 0\\ 0 & 1 & 1\\ 0 & 0 & 1\end{pmatrix}=
\begin{pmatrix}1 & 1 & 1\\ 0 & 1 & 1\\ 0 & 0 & 1\end{pmatrix}" and


"\begin{pmatrix}1 & 0 & 0\\ 0 & 1 & 1\\ 0 & 0 & 1\end{pmatrix}\cdot \begin{pmatrix}1 & 1 & 0\\ 0 & 1 & 0\\ 0 & 0 & 1\end{pmatrix}=
\begin{pmatrix}1 & 1 & 0\\ 0 & 1 & 1\\ 0 & 0 & 1\end{pmatrix}," we conclude that the group "A" is not Abelian.